Suppose you have a credit card with a balance of $14,340.87 that charges 10.13% APR on unpaid balances. The minimum payment each month is $35. If you make just the minimum payment, how much interest will be added to your account balance?
Answer #1

Credit Card Balance = $14340.87

Minimum Payment made = $35

Balance Amount = 14340.87 - 35 = $14305.87

Monthly Interest rate = r = 0.1013/12

Hence, Interest added to account balance = Balance Amount * Interest Rate = 14305.87*0.1013/12 = $120.77